# Self-Evolving Curriculum for LLM Reasoning

### v1

@@ 2.1 Curriculum Selection as Multi-Armed Bandit
This POMDP formulation closely resembles a non-stationary MAB problem, where each arm represents a problem category $c_{i}$, and the objective is to learn the expected return $Q_{t}(c)$ associated with selecting category $c$ at training step $t$. Importantly, the MAB in this context is non-stationary: the expected reward distribution for each arm shifts as the LLM policy is updated over the course of training. To address this well-studied non-stationary bandit problem [51, 49], we leverage the classic TD(0) method [48] to iteratively update $Q_{t}(c)$:
@@ 4 Related Works
Concurrently, two noteworthy studies have explored automatic curriculum learning for RL fine-tuning: Bae et al. [3] propose online filtering of training problems by repeatedly generating solutions to estimate their difficulty, incurring substantial computational overhead; AdaRFT [46] adaptively adjusts curriculum difficulty based on the model’s recent reward signals but relies on explicit difficulty-level ordering, limiting its generalizability to other curriculum categories.
In contrast to these methods, our SEC approach leverages a general, non-stationary MAB formulation to dynamically adjust the curriculum, making it broadly applicable across diverse curriculum categories, arbitrary reward scales, and various RL algorithms.

### v2

@@ 2.1 Curriculum Selection as Multi-Armed Bandit
This POMDP formulation naturally resembles a non-stationary MAB problem, a connection also highlighted by Matiisen et al. [31], where each arm represents a problem category $c_{i}$, and the objective is to learn the expected return $Q_{t}(c)$ associated with selecting category $c$ at training step $t$. Importantly, the MAB in this context is non-stationary: the expected reward distribution for each arm shifts as the LLM policy is updated over the course of training. To address this well-studied non-stationary bandit problem [51, 49], we leverage the classic TD(0) method [48] to iteratively update $Q_{t}(c)$:
@@ 4 Related Works
Concurrently, several studies have explored automatic curriculum learning for RL fine-tuning. Bae et al. [3] propose online filtering of training problems by repeatedly generating solutions to estimate their difficulty. AdaRFT [46] adaptively adjusts curriculum difficulty based on the model’s recent reward signals but relies on explicit difficulty-level ordering. In contrast, SEC leverages a general MAB formulation to dynamically adjust the curriculum. DUMP [56], in parallel to us, also leverages absolute advantage as the curriculum reward with an MAB framework, focusing on the Knights and Knaves logical reasoning puzzle with GRPO. In contrast, our study covers multiple reasoning domains, examines a multi-task RL setting, and validates performance across various RL algorithms.